# Gravitational reheating in Starobinsky inflation

### v2

@@ 2 Inflation and reheating
where $R$ is the Ricci Scalar, $m_{\phi}$ is a mass term, and $M_{\text{P}}=1/\sqrt{8\pi G}$ is the reduced Planck mass. This model fits the observed data from CMB extremely well [22], and is a natural candidate for inflation in that it is a minimal extension of General Relativity, with the inclusion of an additional $R^{2}$ term which is expected to emerge from quantum corrections [23, 24] [^1] One-loop corrections to gravity induce another term $R_{\mu\nu}R^{\mu\nu}$ as well, which are however not part of the standard Starobinsky model which is of interest to us here. An analysis of inflationary dynamics in an effective theory of gravity, including other terms expected to be induced by loop corrections, can be found in refs. [25, 26].. It can be shown [3] that, after performing a conformal transformation on the metric,
@@
Figure 1: Starobinsky potential. The black dashed vertical line divides the inflationary regime and reheating. $\phi_{e}$ represent the inflaton at the end of inflation, when it leaves the plateaux of the potential.
@@
$A_{s}=2.42\times 10^{-9}\quad\text{and}\quad n_{s}=0.9649\qquad(k_{*}=0.002\text{ Mpc}^{-1}).$ (2.14)
@@ 4.1 Bogoliubov transformations and particle creation from vacuum
In addition, the set of modes $\left\{(2\pi)^{-3/2}\chi_{k}(t)e^{-ik\cdot x}\right\}$ must form an orthonormal basis for a space of solutions with positive norm of the Klein-Gordon equation [32]. In order to remove the friction term present in the equation, we introduce the conformal time $d\eta=adt$ and the new rescaled field $\overline{\chi}=a\chi$. This leads to an equation of motion for this new field of the form
@@ 4.2 Numerical method
$\mathcal{H}_{e}=\frac{a^{\prime}(\eta_{e})}{a(\eta_{e})}\simeq 8.8\times 10^{14}~{}\text{GeV},$ (4.10)
@@
Figure 6: Hubble parameter as a function of the number of $e$-folds. The horizontal dashed line represents the value $\mathcal{H}_{e}$ of the conformal Hubble parameter when the first slow-roll parameter satisfies $\epsilon=1$, i.e. at the end of slow-roll inflation, and corresponds to $\mathcal{H}_{e}=8.8\times 10^{14}$ GeV.
@@ 4.3 Case $\mathcal{O}(100~{}\text{GeV})\ll m_{\chi}\ll H_{e}$ and $\xi<1$
Figure 9: (Left) Particle production for $\xi=0$ and $m_{\chi}=0.01H_{e}$. The gray region contains the $k>H_{e}$ modes, where production occurs only in the oscillatory regime. (Right) Evolution of inflaton energy densities and radiation after nonperturbative particle production.
For $\xi=0$ and $m_{\chi}=0.01H_{e}$ this yields the plot shown in figure 9. Notice that the largest contribution to particle production comes from superhorizon modes, $k<H_{e}$. In this window, production takes place even before the oscillations start, in the transition between the slow-roll regime and reheating. There will be some particle production also for subhorizon modes, due to the oscillations of the Ricci scalar, as can be seen in the zoomed in region in figure 9. It can be shown that the amount of produced particles in this window $k>H_{e}$ is the same whether we compute it with the Boltzmann or with the Bogoliubov approaches [36, 18]. However, the contribution from these modes is clearly subdominant. This is just the well-known result that the explosive production in preheating (i.e. in our non-perturbative regime) is much more efficient than the subsequent decay of the inflaton computed perturbatively. Integrating over the entire $k$ range shown in figure 9 (left), we find for the energy density and the corresponding maximum temperature:
$\rho_{\chi}=2.7\times 10^{52}~{}\text{GeV}^{4}\implies T_{\text{MAX}}=5.4\times 10^{12}~{}\text{GeV}.$ (4.21)
@@ 4.4 Case $m_{\chi}\sim\mathcal{O}(100~{}\text{GeV})$
Figure 11: Particle production for $m_{\chi}\lll H_{e}$ considering only modes $k>H_{e}$ where there is no production due to instability.
@@ 4.5 Case $\xi\gg 1$
The enhancement in particle production implies a much larger energy density $\rho_{\chi}$ at the end of reheating. We show in figure 13 the reheaton energy density $\rho_{\chi}$ after 10 $e$-folds as a function of the coupling constant $\xi$ for $m_{\chi}=0.01H_{e}$ and $m_{\chi}=0.001H_{e}$.
@@
Looking at the figure 13, it can be seen that we can have a universe dominated by radiation through the nonperturbative production of particles. However, we cannot define the temperature obtained as the reheating temperature, because after the nonperturbative production, the radiation will decay faster than the inflaton. Then the inflaton dominates again, until the trilinear interactions become relevant. This behavior can be seen in figure 14, where we use as an initial condition the energy densities after the nonperturbative regime for the case where $m_{\chi}=0.01H_{e}$ and $\xi=45$. As $m_{\chi}$ decreases, the density of particles produced grows (as expected from phase space availability), but the energy density decreases (as shown in figure 13) because the energy contribution of each particle decreases by an even larger factor. However, one cannot extrapolate this trend for ever decreasing values of $m_{\chi}$, because eventually the regime of tachyonic instabilities will dominate the production, leading to a much more efficient reheating than the ones we consider in figure 13.
Figure 14: Evolution of inflaton energy densities and radiation after nonperturbative particle production. For $m_{\chi}=0.01H_{e}$ and $\xi=45$.
@@ 5 Conclusions
We found that for $\xi=0$, most of the production occurs at the beginning of reheating, when the field starts rolling down the potential well, and before it starts oscillating around the minimum. The contributions from the oscillatory regime affect mainly the subhorizon modes, and we expect that in this case they are well described by the perturbative approach, as shown in refs. [36, 18]. Thus, our finding that the subhorizon ($k>H_{e}$) contributions are subdominant (cf. figure 9) reproduces the well-known result that the explosive period of preheating is much more efficient for particle production than subsequent inflaton decays [38]. In this case, the maximum reheating temperature after nonperturbative particle production was determined, with the value of $T_{\text{MAX}}=5.4\times 10^{12}~{}\text{GeV}$. After the preheating period, the densities obtained were used as the initial condition for a subsequent perturbative particle production regime. We saw that preheating did not alter the reheating temperature obtained in the perturbative approach.
We then analyzed the case of $\xi>1$. For large $\xi$, the effect of the oscillations of the scalar of curvature in the evolution of the modes of the reheaton is amplified, leading to stronger resonance effects in the $k>H_{e}$ region, enhancing the production of particles. As a result, there is a significant increase in the reheaton energy density, leading to an increase in the maximum temperature achieved. We have seen that for the case where $\xi>45$ and $m_{\chi}=0.01H_{e}\sim 10^{12}$ GeV the universe can become dominated by radiation during preheating. This does not mean that the $\Lambda$CDM is recovered at this point, because radiation may subsequently decay faster than the inflaton, and the latter may dominate again. However, we have shown that the trilinear coupling between the inflaton and the reheaton field will cause a decay of the latter into the former, and, taking this into account, the radiation density does become dominant again once and for all, as in [30]. This is the moment when reheating is complete! Morever, because the trilinear becomes relevant only at the end of reheating, when the perturbative approach is already adequate for a description of inflaton decays, we find that the final reheating temperature is not too different when comparing the perturbative and non-perturbative approaches.

### v3

@@ 2 Inflation and reheating
where $R$ is the Ricci Scalar, $m_{\phi}$ is a mass term, and $M_{\text{P}}=1/\sqrt{8\pi G}$ is the reduced Planck mass. This model fits the observed data from CMB extremely well [22], and is a natural candidate for inflation in that it is a minimal extension of General Relativity, with the inclusion of an additional $R^{2}$ term which is expected to emerge from quantum corrections [23, 24] [^1] One-loop corrections to gravity induce another term $R_{\mu\nu}R^{\mu\nu}$ as well, which are however not part of the standard Starobinsky model which is of interest to us here. An analysis of inflationary dynamics in an effective theory of gravity, including other terms expected to be induced by loop corrections, can be found in refs. [25, 26, 27].. It can be shown [3] that, after performing a conformal transformation on the metric,
@@
Figure 1: Starobinsky potential.
@@
$A_{s}=2.35\times 10^{-9}\quad\text{and}\quad n_{s}=0.9649\qquad(k_{*}=0.002\text{ Mpc}^{-1}).$ (2.14)
@@ 4.1 Bogoliubov transformations and particle creation from vacuum
In addition, the set of modes $\left\{(2\pi)^{-3/2}\chi_{k}(t)e^{-ik\cdot x}\right\}$ must form an orthonormal basis for a space of solutions with positive norm of the Klein-Gordon equation [33]. In order to remove the friction term present in the equation, we introduce the conformal time $ad\eta=dt$ and the new rescaled field $\overline{\chi}=a\chi$. This leads to an equation of motion for this new field of the form
@@ 4.2 Numerical method
$H_{e}=\frac{a^{\prime}(\eta_{e})}{a(\eta_{e})^{2}}\simeq 7.8\times 10^{12}~{}\text{GeV},$ (4.10)
@@
Figure 6: Hubble parameter as a function of the number of $e$-folds. The vertical dashed line represents the $e$-folds when the first slow-roll parameter satisfies $\epsilon=1$, i.e. at the end of slow-roll inflation.
@@ 4.3 Case $\mathcal{O}(100~{}\text{GeV})\ll m_{\chi}\ll H_{e}$ and $\xi<1$
Figure 9: (Left) Particle production for $\xi=0$ and $m_{\chi}=H_{e}$. (Right) Evolution of inflaton energy densities and radiation after nonperturbative particle production.
For $\xi=0$ and $m_{\chi}=H_{e}$ this yields the plot shown in figure 9. Notice that there is particle production for superhorizon modes, $k<H_{e}$ and also for subhorizon modes $k>H_{e}$. It can be shown that the number of produced particles in this window $k>H_{e}$ is the same whether we compute it with the Boltzmann or with the Bogoliubov approaches [37, 18]. Integrating over the entire $k$ range shown in figure 9 (left), we find for the energy density and the corresponding maximum temperature:
$\rho_{\chi}=4.3\times 10^{52}~{}\text{GeV}^{4}\implies T_{\text{MAX}}=6\times 10^{12}~{}\text{GeV}.$ (4.21)
@@ 4.4 Case $m_{\chi}\sim\mathcal{O}(100~{}\text{GeV})$
Figure 11: Particle production for $m_{\chi}\lll H_{e}$ considering only modes $k\gg H_{e}$ where there is no production due to instability.
@@ 4.5 Case $\xi\gg 1$
The enhancement in particle production implies a much larger energy density $\rho_{\chi}$ at the end of reheating. We show in figure 13 the reheaton energy density $\rho_{\chi}$ after 10 $e$-folds as a function of the coupling constant $\xi$ for $m_{\chi}=H_{e}$ and $m_{\chi}=0.1H_{e}$.
@@
Looking at the figure 13, it can be seen that we can have a universe dominated by radiation through the nonperturbative production of particles. However, we cannot define the temperature obtained as the reheating temperature, because after the nonperturbative production, the radiation will decay faster than the inflaton. Then the inflaton dominates again, until the trilinear interactions become relevant. This behavior can be seen in figure 14, where we use as an initial condition the energy densities after the nonperturbative regime for the case where $m_{\chi}=H_{e}$ and $\xi=45$. As $m_{\chi}$ decreases, the density of particles produced grows (as expected from phase space availability), but the energy density decreases (as shown in figure 13) because the energy contribution of each particle decreases by an even larger factor. However, one cannot extrapolate this trend for ever decreasing values of $m_{\chi}$, because eventually the regime of tachyonic instabilities will dominate the production, leading to a much more efficient reheating than the ones we consider in figure 13.
Figure 14: Evolution of inflaton energy densities and radiation after nonperturbative particle production. For $m_{\chi}=H_{e}$ and $\xi=45$.
@@ 5 Conclusions
We found that for $\xi=0$, most of the production occurs at the beginning of reheating, when the field starts rolling down the potential well, and before it starts oscillating around the minimum. This particle production impacts both subhorizon and superhorizon modes. For subhorizon modes, we expect them to be well described by the perturbative approach, as shown in refs. [37, 18]. In this case, the maximum reheating temperature after nonperturbative particle production was determined, with the value of $T_{\text{MAX}}=6\times 10^{12}~{}\text{GeV}$. After the preheating period, the densities obtained were used as the initial condition for a subsequent perturbative particle production regime. We saw that preheating did not alter the reheating temperature obtained in the perturbative approach.
We then analyzed the case of $\xi>1$. For large $\xi$, the effect of the oscillations of the scalar of curvature in the evolution of the modes of the reheaton is amplified, leading to stronger resonance effects in the $k>H_{e}$ region, enhancing the production of particles. As a result, there is a significant increase in the reheaton energy density, leading to an increase in the maximum temperature achieved. We have seen that for the case where $\xi>45$ and $m_{\chi}=H_{e}\sim 10^{12}$ GeV the universe can become dominated by radiation during preheating. This does not mean that the $\Lambda$CDM is recovered at this point, because radiation may subsequently decay faster than the inflaton, and the latter may dominate again. However, we have shown that the trilinear coupling between the inflaton and the reheaton field will cause a decay of the latter into the former, and, taking this into account, the radiation density does become dominant again once and for all, as in [31]. This is the moment when reheating is complete! Morever, because the trilinear becomes relevant only at the end of reheating, when the perturbative approach is already adequate for a description of inflaton decays, we find that the final reheating temperature is not too different when comparing the perturbative and non-perturbative approaches.
@@
Acknowledgements
LM would like to thank CAPES for financial support during the preparation of this work. The authors would also like to thank an anonymous referee for constructive comments and suggestions that have improved the quality of the work.